Could you please help me with this proble
Find the equation of a circle with center at the origin and point (4,7) on the circle.

Answer by stanbon(75887)

The radius is the distance from (0,0) to (4,7)
=sqrt(4^2+7^2)= sqrt(65)
EQUATION:
x^2+y^2=r^2
x^2+y^2=65
Cheers,
Stan H.